OpenStudy (anonymous):

Write the function in slope-intercept form. Then graph the function. 2x - 5y = -6

OpenStudy (anonymous):

first, solve for y (get y by itself)

OpenStudy (shamil98):

y = mx + b is your slope intercept form. 2x -5y = -6 -2x -2x -5y = -2x -6 divide everything by -5 y = 2/5 x + 6/5

OpenStudy (shamil98):

6/5 is equal to 1.2 , so your y-intercept would be (0,1.2) And you graph the equation from there with (up2/right5) and (down2/left5)